# SUMMA: A Multimodal Large Language Model for Advertisement Summarization

### v1

@@ Abstract.
Understanding multimodal video ads is crucial for improving query-ad matching and relevance ranking on short video platforms, enhancing advertising effectiveness and user experience. However, the effective utilization of multimodal information with high commercial value still largely constrained by reliance on highly compressed video embeddings—has long been inadequate. To address this, we propose SUMMA (the abbreviation of SUmmarizing MultiModal Ads), a multimodal model that automatically processes video ads into summaries highlighting the content of highest commercial value, thus improving their comprehension and ranking in Douyin search-advertising systems. SUMMA is developed via a two-stage training strategy—multimodal supervised fine-tuning followed by reinforcement learning with a mixed reward mechanism—on domain-specific data containing video frames and ASR/OCR transcripts, generating commercially valuable and explainable summaries. We integrate SUMMA-generated summaries into our production pipeline, directly enhancing the candidate retrieval and relevance ranking stages in real search-advertising systems. Both offline and online experiments show substantial improvements over baselines, with online results indicating a statistically significant 1.5% increase in advertising revenue. Our work establishes a novel paradigm for condensing multimodal information into representative texts, effectively aligning visual ad content with user query intent in retrieval and recommendation scenarios.
@@ 1. Introduction
In this paper, we propose SUMMA (SUmmarizing MultiModal Ads), a specialized MLLM for search advertising applications. Initially, we perform supervised fine-tuning (SFT) from Qwen2.5-VL-3B-Instruct on an extensive collection of advertising short video summarization data (sourced from Douyin). To ensure the quality of our SFT data, we implement a rule-based filtering approach leveraging LLM-as-a-judge. This methodology employs dual criteria: assessing the intrinsic quality of video summaries and evaluating their potential to enhance performance in downstream relevance tasks. By concatenating advertisement videos and their corresponding OCR/ASR texts as the input, the yielded SUMMA-SFT model exhibits a strong basic capacity for summarizing multimodal advertisements. Subsequently, we further enhance SUMMA-SFT through reinforcement learning using high-quality meticulously annotated data. Our RL reward design draws inspiration from (Feng et al., 2025a) and (Chang et al., 2025), where the model receives encouragement or penalty signals based on various metrics with a reference (e.g., BLEU and external LLM judgment), depending on the generated outputs and human-annotated ground truth responses.
@@ 3. Methodology
In this section, we present SUMMA, an MLLM that generates representative, concise, and commercially oriented summaries by integrating advertising videos with their corresponding OCR/ASR transcripts. As a roadmap for what follows, Figure 2 illustrates the overall architecture of SUMMA. Building on the Qwen2.5-VL, the framework jointly encodes videos as well as the corresponding OCR/ASR transcripts, then decodes them into advertising-style highlights that emphasize product brands. To optimize this ability, we adopt a two-stage pipeline. 1. Supervised Fine-Tuning (SFT) Stage:We perform data collection and processing by gathering advertising videos together with their OCR/ASR transcripts. Employing cold-start multimodal supervised fine-tuning with data generated and cleaned by Doubao-1.5-pro (ByteDance, 2025) (hereafter abbreviated as Doubao). 2. Reinforcement Learning (RL) Stage: Utilizing GRPO combined with a mixed-reward mechanism to further align the model with human preferences. The remainder of this section details each component of the pipeline in turn.
@@ 3.1. Data Curation
Leveraging our in-house search-ad video logs, we curate three complementary resources—AdSum-Doubao (LLM-bootstrapped summaries for supervised fine-tuning, SFT), AdSum-Human (expert-annotated summaries for reinforcement learning, RL), and AdSum-Test (a held-out set reserved for final evaluation). The construction of each dataset is detailed in the following sections.
@@
To evaluate the ad video summarization capabilities of our SUMMA, we collect 2.5k samples from Douyin for manual annotation using the aforementioned methodology.
@@ 4.2. Experiment Setup
For SFT, we take Qwen2.5-VL as our base model to train SUMMA-SFT, with a learning rate of 5e-7. For learning rate scheduler, we adopt a cosine decay strategy and a linear warm-up of 2000 steps. We use the AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999, and a weight decay rate of 0.02.
@@ 4.3. Offline Experiments
  • Qwen2.5-VL different variants (3B and 7B Instruct versions)
@@
Qwen2.5-VL-3B-Instruct 0.10 0.156 0.26 0.09
Qwen2.5-VL-7B-Instruct 0.14 0.18 0.29 0.12
@@
As shown in Table 2, our method demonstrates consistent improvements across all evaluation metrics. Even our intermediary SUMMA-SFT model can outperform both Qwen2.5-VL models by significant margins (e.g., 0.36 vs 0.14 in BLEU, and 1.01 vs 0.18 in CIDEr). This performance gap highlights the effectiveness of our constructed datasets, which infuse domain-specific knowledge into the model and thus enhance its capability in search advertising scenarios. Moreover, compared to the other training strategies (two-stage SFT or mixed 1-stage SFT), “SFT then RL” is found to be the best one, providing additional evidence that the training paradigm also remains effective when applied to the search advertising domain, and thus demonstrating its broader applicability. Figure 8 is a case study showing the superiority of the summary generated from SUMMA over the base model.
@@ 4.4. Online Experiments
A production-level A/B testing is implemented by incorporating our SUMMA into the BERT-based retrieval and relevance models within our real-time search advertising system. While maintaining all other factors unchanged, we expose the experimental variant to 10% of total search ad traffic of Douyin for 14 consecutive days to ensure statistical significance.

### v2

@@ Abstract.
Understanding multimodal video ads is crucial for improving query-ad matching and relevance ranking on short video platforms, enhancing advertising effectiveness and user experience. However, the effective utilization of multimodal information with high commercial value still largely constrained by reliance on highly compressed video embeddings—has long been inadequate. To address this, we propose SUMMA (the abbreviation of SUmmarizing MultiModal Ads), a multimodal model that automatically processes video ads into summaries highlighting the content of highest commercial value, thus improving their comprehension and ranking in our search-advertising systems. SUMMA is developed via a two-stage training strategy—multimodal supervised fine-tuning followed by reinforcement learning with a mixed reward mechanism—on domain-specific data containing video frames and ASR/OCR transcripts, generating commercially valuable and explainable summaries. We integrate SUMMA-generated summaries into our production pipeline, directly enhancing the candidate retrieval and relevance ranking stages in real search-advertising systems. Both offline and online experiments show substantial improvements over baselines, with online results indicating a statistically significant 1.5% increase in advertising revenue. Our work establishes a novel paradigm for condensing multimodal information into representative texts, effectively aligning visual ad content with user query intent in retrieval and recommendation scenarios.
@@ 1. Introduction
In this paper, we propose SUMMA (SUmmarizing MultiModal Ads), a specialized MLLM for search advertising applications. Initially, we perform supervised fine-tuning (SFT) from Qwen2-VL-2B-Instruct on an extensive collection of advertising short video summarization data (sourced from our short video platform). To ensure the quality of our SFT data, we implement a rule-based filtering approach leveraging LLM-as-a-judge. This methodology employs dual criteria: assessing the intrinsic quality of video summaries and evaluating their potential to enhance performance in downstream relevance tasks. By concatenating advertisement videos and their corresponding OCR/ASR texts as the input, the yielded SUMMA-SFT model exhibits a strong basic capacity for summarizing multimodal advertisements. Subsequently, we further enhance SUMMA-SFT through reinforcement learning using high-quality meticulously annotated data. Our RL reward design draws inspiration from (Feng et al., 2025a) and (Chang et al., 2025), where the model receives encouragement or penalty signals based on various metrics with a reference (e.g., BLEU and external LLM judgment), depending on the generated outputs and human-annotated ground truth responses.
@@ 3. Methodology
In this section, we present SUMMA, an MLLM that generates representative, concise, and commercially oriented summaries by integrating advertising videos with their corresponding OCR/ASR transcripts. As a roadmap for what follows, Figure 2 illustrates the overall architecture of SUMMA. Building on the Qwen2-VL, the framework jointly encodes videos as well as the corresponding OCR/ASR transcripts, then decodes them into advertising-style highlights that emphasize product brands. To optimize this ability, we adopt a two-stage pipeline. 1. Supervised Fine-Tuning (SFT) Stage:We perform data collection and processing by gathering advertising videos together with their OCR/ASR transcripts. Employing cold-start multimodal supervised fine-tuning with data generated and cleaned by Doubao-1.5-pro (ByteDance, 2025) (hereafter abbreviated as Doubao). 2. Reinforcement Learning (RL) Stage: Utilizing GRPO combined with a mixed-reward mechanism to further align the model with human preferences. The remainder of this section details each component of the pipeline in turn.
@@ 3.1. Data Curation
Leveraging our in-house search-ad video data, we curate three complementary resources—AdSum-Doubao (LLM-bootstrapped summaries for supervised fine-tuning, SFT), AdSum-Human (expert-annotated summaries for reinforcement learning, RL), and AdSum-Test (a held-out set reserved for final evaluation). The construction of each dataset is detailed in the following sections.
@@
To evaluate the ad video summarization capabilities of our SUMMA, we collect 2.5k samples from our platform for manual annotation using the aforementioned methodology.
@@ 4.2. Experiment Setup
For SFT, we take Qwen2-VL as our base model to train SUMMA-SFT, with a learning rate of 5e-7. For learning rate scheduler, we adopt a cosine decay strategy and a linear warm-up of 2000 steps. We use the AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999, and a weight decay rate of 0.02.
@@ 4.3. Offline Experiments
  • Qwen2-VL different variants (2B and 7B Instruct versions)
@@
Qwen2-VL-2B-Instruct 0.10 0.156 0.26 0.09
Qwen2-VL-7B-Instruct 0.14 0.18 0.29 0.12
@@
As shown in Table 2, our method demonstrates consistent improvements across all evaluation metrics. Even our intermediary SUMMA-SFT model can outperform both Qwen2-VL models by significant margins (e.g., 0.36 vs 0.14 in BLEU, and 1.01 vs 0.18 in CIDEr). This performance gap highlights the effectiveness of our constructed datasets, which infuse domain-specific knowledge into the model and thus enhance its capability in search advertising scenarios. Moreover, compared to the other training strategies (two-stage SFT or mixed 1-stage SFT), “SFT then RL” is found to be the best one, providing additional evidence that the training paradigm also remains effective when applied to the search advertising domain, and thus demonstrating its broader applicability. Figure 8 is a case study showing the superiority of the summary generated from SUMMA over the base model.
@@ 4.4. Online Experiments
A production-level A/B testing is implemented by incorporating our SUMMA into the BERT-based retrieval and relevance models within our real-time search advertising system. While maintaining all other factors unchanged, we expose the experimental variant to 10% of total search ad traffic of our platform for 14 consecutive days to ensure statistical significance.